### 16 Mar 2026 v1

# Acceleration of multi-component multiple-precision arithmetic with branch-free algorithms and SIMD vectorization

Abstract
Multiple-precision floating-point branch-free algorithms can significantly accelerate multi-component arithmetic implemented by combining hardware-based binary64 and binary32, particularly for triple- and quadruple-precision computations. In this study, we achieved benchmark results on x86 and ARM CPU platforms to quantify the accelerations achieved in linear computations and polynomial evaluation by integrating these algorithms.
1 Introduction
Recent advances in artificial intelligence (AI), underpinned by deep learning technologies, have led to the widespread adoption of large language models and generative AI systems, permeating diverse industrial sectors and everyday human activities, including software development and coding. This trend has accelerated the shortening of mantissa length of floating-point formats employed in AI computation in pursuit of higher performance. However, in scientific and engineering computing, double-precision arithmetic represents the minimum standard, and ill-conditioned problems frequently demand even higher precision. In such cases, the effective exploitation of single instruction multiple data (SIMD) instructions and OpenMP-based parallelization techniques is crucial for harnessing the full capabilities of modern hardware. Among existing multiple-precision arithmetic frameworks, arbitrary-precision arithmetic is dominated by MPFR [15], which relies on the multiple-precision natural number arithmetic (MPN) kernel of GMP [5], and MPC [3]—an arbitrary-precision complex arithmetic library built on MPFR that achieves acceleration through the Karatsuba 3M multiplication algorithm—is widely used.
By contrast, fixed-precision multi-component arithmetic based on error-free transformation (EFT) techniques, which employs combinations of multiple IEEE binary32 (24-bit mantissa and 32-bit total length) and binary64 values supported by existing hardware, is represented by quadruple-double (QD) [1] for CPUs and GQD [12]—a CUDA port of QD, though dated—for GPUs. When a relatively short mantissa length suffices, such as in double-double (DD, 106-bit), triple-double (TD, 159-bit), and quadruple-double (QD, 212-bit) arithmetic implemented by combining two to four binary64 values, computations can be executed faster than with MPFR. It has further been demonstrated by Kotakemori et al.’s Lis [7] and Hishinuma et al. [6] that basic linear computations can be accelerated by leveraging AVX, AVX2, and AVX-512, which are x86-based SIMD instruction sets. Building on an existing AVX2 implementation, we have demonstrated that basic linear computations in DD, TD, and QD precision can be accelerated on Raspberry Pi and Snapdragon platforms by exploiting Neon instructions [18].
Nevertheless, DD, TD, and QD arithmetic invariably require a renormalization process at the conclusion of each operation. Particularly, for TD and QD arithmetic, this step has been identified as a bottleneck impeding performance gains through SIMD vectorization. Double-precision pair arithmetic that bypasses renormalization has been proposed as a faithfully rounded pair arithmetic scheme [11], and practical sparse matrix solver implementations that exploit its efficiency have been reported [14]. The branch-free algorithms for multiple-precision floating-point arithmetic addressed in this work [17] eliminate all conditional branches from every algorithmic step, including renormalization, and additionally reduce the operation count for TD and QD arithmetic relative to conventional algorithms. Consequently, this methodology can substantially accelerate triple- and quadruple-precision multi-component arithmetic implemented by combining hardware-based binary64 and binary32 values, and is immediately applicable to existing fixed-precision multiple-precision arithmetic implementations. As the benchmarks presented in the original paper provided insufficient comparison against existing SIMD implementations, we first quantified the speedups achievable in basic linear computations by incorporating these algorithms on x86 and ARM CPU platforms. Furthermore, as a practical application, we analyzed the effectiveness of the proposed approach when applied to a multiple-precision algebraic equation solver.
2 Branch-Free Algorithms for DD, TD, and QD Addition and Multiplication
The extension of fixed-precision floating-point numbers via EFT originated with Dekker [2]. Multiple-precision arithmetic with extended mantissa length is implemented by employing EFT primitives, represented by QuickTwoSum (Algorithms 1, 2, 3), which return both the rounded result and its rounding error using only existing floating-point operations (addition $\oplus$, subtraction $\ominus$, and multiplication $\otimes$). In the following, all variables are assumed to be representable exactly as normalized floating-point numbers.
QuickTwoSum (Algorithm 1) returns the floating-point addition result $s=a\oplus b$ and its rounding error $e$, but only under the precondition $|a|\geq|b|$; provided no overflow occurs, it maintains the relation $a+b=s+e$.
Algorithm 1 ($s$, $e$) :=:= QuickTwoSum($a$, $b$)
$s:=a\oplus b$; $e:=b\ominus(s\ominus a)$; return ($s$, $e$)
TwoSum (Algorithm 2) returns the floating-point addition result $s=a\oplus b$ and its rounding error $e$ unconditionally, regardless of the relative magnitudes of $a$ and $b$; it likewise maintains the relation $a+b=s+e$.
Algorithm 2 ($s$, $e$) :=:= TwoSum($a$, $b$)
$s:=a\oplus b$; $v:=s\ominus a$; $e:=(a\ominus(s\ominus v))\oplus(b\ominus v)$; return($s$, $e$)
TwoProd (Algorithm 3) returns the floating-point multiplication result $p=a\otimes b$ and its rounding error $e$, maintaining the relation $ab=p+e$. As virtually all floating-point hardware currently in operation supports fused multiply-add (FMA) instructions, the FMA-based variant of TwoProd was employed throughout this work.
Algorithm 3 ($p$, $e$) :=:= TwoProd($a$, $b$)
$p:=a\otimes b$; $e:=\mbox{FMA}(a,b,-p)$ ($a\times b-p$); return($p$, $e$)
As these EFT operations are constructed entirely from existing floating-point operations and contain no conditional branches, they can be directly implemented using SIMD instructions. With AVX2, which employs 256-bit registers on x86 architectures, four binary64 or eight binary32 values can be processed simultaneously by EFT operations; with Neon on AArch64, which employs 128-bit registers, two binary64 or four binary32 values can be processed simultaneously. Computations that contain no conditional branches and can be executed without overflowing SIMD registers in this manner are termed branch-free (BF).
By combining EFT operations, multiple-precision arithmetic with an extended mantissa can be implemented by composing multiple floating-point numbers; this approach is referred to as the multi-component (or multiterm) method. Because subtraction reduces to sign negation and division can be implemented via Newton’s method combining addition and multiplication, these two operations are crucial from a performance perspective. The correct handling of rounding errors is critical in this implementation and requires a carefully derived, rigorous error analysis. As the mantissa length increases, the operation count grows, and error management becomes increasingly complex. Zhang & Aiken proposed an abstract model for automated reasoning that simplifies floating-point arithmetic [16] and introduced branch-free multi-component multiple-precision addition and multiplication algorithms. In the following, we present and discuss these BF algorithms.
2.1 Remarks on double-word (DW) Arithmetic
The simplest form of multi-component multiple-precision floating-point arithmetic is DW arithmetic. When binary64 (double-precision) is used as the base format, it is abbreviated as DD; when binary32 (single-precision) is used, it is abbreviated as DS. Algorithm 4 presents the default DW addition currently defined in QD [1], which is the sloppy or fast variant; its final rounding error is known to be larger than that of the accurate version (Algorithm 5).
Algorithm 4 $c[2]:=$ DWAdd($a[2]$, $b[2]$)
$(s,e):=\mbox{TwoSum}(a[0],b[0])$;
$e:=e\oplus(a[1]\oplus b[1])$;
$(c[0],c[1]):=\mbox{QuickTwoSum}(s,e)$;
return($c[2]$)
Algorithm 5 $c[2]:=$ DWAccurateAdd($a[2]$, $b[2]$)
$(s_{1},s_{2}):=\mbox{TwoSum}(a[0],b[0])$;
$(t_{1},t_{2}):=\mbox{TwoSum}(a[1],b[1])$;
$s_{2}:=s_{2}\oplus t_{1}$;
$(s_{1},s_{2}):=\mbox{QuickTwoSum}(s1,s2)$;
$s_{2}:=s_{2}\oplus t_{2}$;
$(c[0],c[1]):=\mbox{QuickTwoSum}(s1,s2)$;
return($c[2]$)
Zhang & Aiken proposed Algorithm 6 as a branch-free reformulation of the accurate DW addition.
Algorithm 6 $c[2]:=$ DWBFAdd($a[2]$, $b[2]$)
$(g_{1},g_{1e}):=\mbox{TwoSum}(a[0],b[0])$;
$(g_{2},g_{2e}):=\mbox{TwoSum}(a[1],b[1])$;
$(g_{3},g_{3e}):=\mbox{QuickTwoSum}(g_{1},g_{2})$;
$g_{4}:=g_{1e}\oplus g_{2e}$;
$g_{5}:=g_{4}\oplus g_{3e}$;
$(c[0],c[1]):=\mbox{QuickTwoSum}(g_{3},g_{5})$;
return($c[2]$)
Algorithm 7 presents the DW multiplication defined as the default in QD.
Algorithm 7 $c[2]:=$ DWMul($a[2]$, $b[2]$)
$(p_{1},p_{2}):=\mbox{TwoProd}(a[0],b[0])$;
$p_{2}:=p_{2}\oplus(a[0]\otimes b[1]\oplus a[1]\otimes b[0])$;
$(c[0],c[1]):=\mbox{QuickTwoProd}(p_{1},p_{2})$;
return($c[2]$)
The corresponding BF multiplication is proposed as presented in Algorithm 8. This algorithm can be interpreted as performing all necessary multiplications, followed by arranging error-handling steps in a branch-free manner.
Algorithm 8 $c[2]:=$ DWBFMul($a[2]$, $b[2]$)
$(p_{00},p_{e00}):=\mbox{TwoProd}(a[0],b[0])$;
$p_{01}:=a[0]\otimes b[1]$;
$p_{10}=a[1]\otimes b[0]$;
$g_{1}:=p_{01}\oplus p_{10}$;
$g_{2}:=p_{e00}\oplus g_{1}$;
$(c[0],c[1]):=\mbox{QuickTwoSum}(p_{00},g_{2})$;
return($c[2]$)
Consequently, the BF variants of DW addition and multiplication involve no fewer floating-point operations than their conventional counterparts, and—as will be discussed subsequently—no measurable performance benefit was observed under the current default DW arithmetic. Therefore, in the benchmark tests reported as follows, DD arithmetic employs the conventional (nonBF) implementations unless otherwise stated; only TD and QD arithmetic utilize the BF algorithms presented hereafter.
2.2 Remarks on TW Arithmetic
For triple-precision arithmetic (Triple-Word, TW), one approach is to consider the fourth component of QD arithmetic as zero; alternatively, Fabiano et al. [4] proposed a formulation with a reduced operation count. Through SIMD implementation, we validated that the former approach outperforms the latter [9]; therefore, Algorithm 9 is used for TW addition and Algorithm 10 for TW multiplication.
Algorithm 9 $c[3]:=$ TWAdd($a[3]$, $b[3]$)
$s_{0}:=a[0]\oplus b[0]$; $s_{1}:=a[1]\oplus b[1]$; $s_{2}:=a[2]\oplus b[2]$;
$v_{0}:=s_{0}\ominus a[0]$$v_{1}:=s_{1}\ominus a[1]$$v_{2}:=s_{2}\ominus a[2]$;
$u_{0}:=s_{0}\ominus v_{0}$; $u_{1}:=s_{1}\ominus v_{1}$; $u_{2}:=s_{2}\ominus v_{2}$;
$w_{0}:=a[0]\ominus u_{0}$$w_{1}:=a[1]\ominus u_{1}$; $w_{2}:=a[2]\ominus u_{2}$;
$u_{0}:=b[0]\ominus v_{0}$; $u_{1}:=b[1]\ominus v_{1}$; $u_{2}:=b[2]\ominus v_{2}$;
$t_{0}:=w_{0}\oplus u_{0}$; $t_{1}:=w_{1}\oplus u_{1}$; $t_{2}:=w_{2}\oplus u_{2}$;
$(s_{1},t_{0}):=\mbox{TwoSum}(s_{1},t_{0})$; $(i_{1},i_{2}):=\mbox{TwoSum}(s_{2},t_{0})$;
$(s_{2},i_{3}):=\mbox{TwoSum}(t_{1},i_{1})$; $(t_{0},t_{1}):=\mbox{TwoSum}(i_{2},i_{3})$ $t_{0}:=t_{0}\oplus t_{1}\oplus t_{2}$;
$c[3]:=\mbox{TWRenormalize}(s_{0},s_{1},s_{2},t_{0})$;
return($c[3]$)
Algorithm 10 $c[3]:=$ TWMul($a[3]$, $b[3]$)
$(p_{0},q_{0}):=\mbox{TwoProd}(a[0],b[0])$; $(p_{1},q_{1}):=\mbox{TwoProd}(a[0],b[1])$;
$(p_{2},q_{2}):=\mbox{TwoProd}(a[1],b[0])$;
$(p_{3},q_{3}):=\mbox{TwoProd}(a[0],b[2])$; $(p_{4},q_{4}):=\mbox{TwoProd}(a[1],b[1])$;
$(p_{5},q_{5}):=\mbox{TwoProd}(a[2],b[0])$;
$(i_{1},i_{2}):=\mbox{TwoSum}(p_{1},p_{2})$; $(p_{1},i_{3}):=\mbox{TwoSum}(q_{0},i_{1})$; $(p_{2},q_{0}):=\mbox{TwoSum}(i_{2},i_{3})$;
$(i_{1},i_{2}):=\mbox{TwoSum}(p_{2},q_{1})$; $(p_{2},i_{3}):=\mbox{TwoSum}(q_{2},i_{1})$; $(q_{1},q_{2}):=\mbox{TwoSum}(i_{2},i_{3})$;
$(i_{1},i_{2}):=\mbox{TwoSum}(p_{3},p_{4})$; $(p_{3},i_{3}):=\mbox{TwoSum}(p_{5},i_{1})$; $(p_{4},p_{5}):=\mbox{TwoSum}(i_{2},i_{3})$;
$(s_{0},t_{0}):=\mbox{TwoSum}(p_{2},p_{3})$; $(s_{1},t_{1}):=\mbox{TwoSum}(q_{1},p_{4})$; $s_{2}:=q_{2}\oplus p_{5}$;
$(s_{1},t_{0}):=\mbox{TwoSum}(s_{1},t_{0})$; $s_{2}:=s_{2}\oplus(t_{0}\oplus t_{1})$; $(q_{0},q_{3}):=\mbox{TwoSum}(q_{0},q_{3})$;
$(q_{4},q_{5}):=\mbox{TwoSum}(q_{4},q_{5})$; $(t_{0},t_{1}):=\mbox{TwoSum}(q_{0},q_{4})$; $t_{1}:=t_{1}\oplus(q_{3}\oplus q_{5})$;
$(t_{0},t_{1}):=\mbox{TwoSum}(q_{3},s_{1})$;
$c[3]:=\mbox{TWRenormalize}(p_{0},p_{1},s_{0},t_{0})$;
return($c[3]$)
Both algorithms conclude with a call to the TW renormalization function TWRenormalize, which requires up to five QuickTwoSum operations and necessitates conditional branching via if-statements. This function has not yet been SIMD-vectorized and is considered a primary source of performance degradation.
In contrast, the proposed BF algorithms, presented by Algorithms 11 and 12, introduce no conditional branches in the renormalization-equivalent steps and require fewer operations than their conventional counterparts. This renders them highly amenable to performance improvement through SIMD vectorization.
Algorithm 11 $c[3]:=$ TWBFAdd($a[3]$, $b[3]$)
$(a_{1},b_{1}):=\mbox{TwoSum}(a[0],b[0])$; $(c_{1},d_{1}):=\mbox{TwoSum}(a[1],b[1])$;
$(e_{1},f_{1}):=\mbox{TwoSum}(a[2],b[2])$;
$(a_{2},c_{2}):=\mbox{QuickTwoSum}(a_{1},c_{1})$; $b_{2}:=b_{1}\oplus f_{1}$;
$(d_{2},e_{2}):=\mbox{TwoSum}(d_{1},e_{1})$; $(a_{3},d_{3}):=\mbox{QuickTwoSum}(a_{2},d_{2})$;
$(b_{3},c_{3}):=\mbox{TwoSum}(b_{2},c_{2})$; $c_{4}:=c_{3}\oplus e_{2}$;
$(c_{5},d_{5}):=\mbox{TwoSum}(c_{4},d_{3})$; $(b_{6},c_{6}):=\mbox{TwoSum}(b_{3},c_{5})$;
$(c[0],b_{7})=\mbox{QuickTwoSum}(a_{3},b_{6})$; $c_{7}:=c_{6}\oplus d_{5}$;
$(c[1],c[2]):=\mbox{QuickTwoSum}(b_{7},c_{7})$;
return($c[3]$)
Algorithm 12 $c[3]:=$ TWBFMul($a[3]$, $b[3]$)
$(a_{0},b_{0}):=\mbox{TwoProd}(a[0],b[0])$; $(c_{0},e_{0}):=\mbox{TwoProd}(a[0],b[1])$;
$(d_{0},f_{0}):=\mbox{TwoProd}(a[1],b[0])$;
$g_{0}:=a[0]\otimes b[2]$; $h_{0}:=a[1]\otimes b[1]$; $i_{0}:=a[2]\otimes b[0]$;
$(c_{1},d_{1}):=\mbox{TwoSum}(c_{0},d_{0})$; $e_{1}:=e_{0}\oplus f_{0}$; $g_{1}:=g_{0}\oplus i_{0}$;
$(b_{2},c_{2}):=\mbox{TwoSum}(b_{0},c_{1})$; $g_{2}:=g_{1}\oplus h_{0}$; $(a_{3},b_{3}):=\mbox{QuickTwoSum}(a_{0},b_{2})$;
$c_{3}:=c_{2}\oplus d_{1}$; $e_{3}:=e_{1}\oplus g_{2}$; $c_{4}:=c_{3}\oplus e_{3}$;
$(b_{5},c_{5}):=\mbox{QuickTwoSum}(b_{3},c_{4})$; $(c[0],b_{6}):=\mbox{QuickTwoSum}(a_{3},b_{5})$;
$(c[1],c[2]):=\mbox{QuickTwoSum}(b_{6},c_{5})$;
return($c[3]$)
2.3 Remarks on QW Arithmetic
Quadruple-precision arithmetic (Quadruple-Word, QW) is based on the implementation provided in QD; as noted previously, TD arithmetic is an abbreviated form of QD arithmetic with the fourth component zeroed. Both QD addition and QD multiplication conclude with a call to the QW renormalization function (QWRenormalize), which contains conditional branches and requires up to seven QuickTwoSum function calls.
In contrast, the BF algorithms are presented by Algorithms 13 and 14. Although the time required to discover these BF algorithms is unknown, the resulting algorithms are remarkably concise and involve a substantially reduced operation count. Consequently, the benefit of SIMD vectorization is expected to be even more pronounced for QW arithmetic than for TW arithmetic.
Algorithm 13 $c[4]:=$ QWBFAdd($a[4]$, $b[4]$)
$(a_{1},b_{1}):=\mbox{TwoSum}(a[0],b[0])$; $(c_{1},d_{1}):=\mbox{TwoSum}(a[1],b[1])$;
$(e_{1},f_{1}):=\mbox{TwoSum}(a[2],b[2])$;
$(g_{1},h_{1}):=\mbox{TwoSum}(a[3],b[3])$; $(a_{2},c_{2}):=\mbox{QuickTwoSum}(a_{1},c_{1})$;
$b_{2}:=b_{1}\oplus h_{1}$; $(d_{2},e_{2}):=\mbox{TwoSum}(d_{1},e_{1})$;
$(f_{2},g_{2}):=\mbox{TwoSum}(f_{1},g_{1})$; $(b_{3},g_{3}):=\mbox{TwoSum}(b_{2},g_{2})$;
$(c_{3},d_{3}):=\mbox{QuickTwoSum}(c_{2},d_{2})$; $(e_{3},f_{3}):=\mbox{TwoSum}(e_{2},f_{2})$;
$(a_{4},c_{4}):=\mbox{QuickTwoSum}(a_{2},c_{3})$; $(d_{4},e_{4}):=\mbox{QuickTwoSum}(d_{3},e_{3})$;
$(b_{5},d_{5}):=\mbox{TwoSum}(b_{3},d_{4})$; $e_{5}:=e_{4}\oplus f_{3}$;
$(b_{6},c_{6}):=\mbox{TwoSum}(b_{5},c_{4})$; $(d_{6},e_{6}):=\mbox{TwoSum}(d_{5},e_{5})$;
$(a_{7},b_{7}):=\mbox{QuickTwoSum}(a_{4},b_{6})$; $(c_{7},d_{7}):=\mbox{QuickTwoSum}(c_{6},d_{6})$;
$e_{8}:=e_{6}\oplus g_{3}$; $(b_{8},c_{8}):=\mbox{QuickTwoSum}(b_{7},c_{7})$;
$d_{9}:=d_{7}\oplus e_{8}$; $(c[0],b_{10}):=\mbox{QuickTwoSum}(a_{7},b_{8})$;
$(c_{10},d_{10}):=\mbox{QuickTwoSum}(c_{8},d_{9})$; $(c[1],c_{11}):=\mbox{QuickTwoSum}(b_{10},c_{10})$;
$(c[2],c[3]):=\mbox{QuickTwoSum}(c_{11},d_{10})$;
return($c[4]$)
Algorithm 14 $c[4]:=$ QDBFMul($a[4]$, $b[4]$)
  $(a_{0},b_{0}):=\mbox{TwoProd}(a[0],b[0])$; $(c_{0},e_{0}):=\mbox{TwoProd}(a[0],b[1])$;
  $(d_{0},f_{0}):=\mbox{TwoProd}(a[1],b[0])$; $(g_{0},j_{0}):=\mbox{TwoProd}(a[0],b[2])$;
  $(h_{0},k_{0}):=\mbox{TwoProd}(a[1],b[1])$ $(i_{0},l_{0}):=\mbox{TwoProd}(a[2],b[0])$;
  $m_{0}:=a[0]\otimes b[3]$; $n_{0}:=a[1]\otimes b[2]$; $o_{0}:=a[2]\otimes b[1]$;
  $p_{0}:=a[3]\otimes b[0]$; $(c_{1},d_{1}):=\mbox{TwoSum}(c_{0},d_{0})$;
  $(e_{1},f_{1}):=\mbox{TwoSum}(e_{0},f_{0})$; $(g_{1},i_{1}):=\mbox{TwoSum}(g_{0},i_{0})$;
  $j_{1}:=j_{0}\oplus l_{0}$; $m_{1}:=m_{0}\oplus p_{0}$; $n_{1}:=n_{0}\oplus o_{0}$;
  $(b_{2},c_{2}):=\mbox{TwoSum}(b_{0},c_{1})$; $(e_{2},h_{2}):=\mbox{TwoSum}(e_{1},h_{0})$;
  $f_{2}:=f_{1}\oplus j_{1}$; $i_{2}:=i_{1}\oplus k_{0}$; $m_{2}:=m_{1}\oplus n_{1}$;
  $(a_{3},b_{3}):=\mbox{QuickTwoSum}(a_{0},b_{2})$; $(c_{3},d_{3}):=\mbox{QuickTwoSum}(c_{2},d_{1})$;
  $(e_{3},g_{3}):=\mbox{TwoSum}(e_{2},g_{1})$; $f_{3}:=f_{2}\oplus m_{2}$; $h_{3}:=h_{2}\oplus i_{2}$;
  $(c_{4},e_{4}):=\mbox{TwoSum}(c_{3},e_{3})$; $d_{4}:=d_{3}\oplus h_{3}$; $f_{4}:=f_{3}\oplus g_{3}$;
  $d_{5}:=d_{4}\oplus e_{4}$; $(c_{6},d_{6}):=\mbox{TwoSum}(c_{4},d_{5})$; $(b_{7},c_{7}):=\mbox{TwoSum}(b_{3},c_{6})$;
  $d_{7}:=d_{6}\oplus f_{4}$; $(c[0],b_{8}):=\mbox{QuickTwoSum}(a_{3},b_{7})$;
  $(c_{8},d_{8}):=\mbox{TwoSum}(c_{7},d_{7})$; $(c[1],c_{9}):=\mbox{TwoSum}(b_{8},c_{8})$;
  $(c[2],c[3]):=\mbox{QuickTwoSum}(c_{9},d_{8})$;
  return($c[4]$)
3 Benchmark Tests for Real and Complex Square Matrix Multiplication
As discussed previously, the adoption of BF arithmetic is expected to yield significant performance gains through SIMD vectorization for triple- and quadruple-precision computations. Here, we report performance evaluation results for real and complex square matrix multiplication obtained by integrating BF arithmetic into our multiple-precision basic linear computation library BNCmatmul [8], executed on the following x86 AVX2 and Arm Neon environments.
EPYC
AMD EPYC 9354P 3.7 GHz 32 cores, Ubuntu 20.04.6 LTS, Intel Compiler version 2021.10.0, GNU MP 6.2.1, MPFR 4.1.0, MPC 1.2.1
Snapdragon
Lenovo IdeaPad Slim 5(Snapdragon X Plus X1P-42-100, 3.4 GHz, 34 MB L3 cache, 8 cores)Windows 11, WSL2 + Ubuntu 24.04.2 LTS, GCC 13.3.0
Standard intrinsic functions provided as C functions were used for SIMD implementation in both environments. The Arm Neon vectorization and OpenMP parallelization were implemented primarily with the assistance of Anthropic Claude, by converting the existing AVX2 code [18].
3.1 Real Matrix Multiplication
The square matrices employed in the tests are defined as
$\begin{split}C&:=AB,\\ A&=[a_{ij}]_{i,j=1,2,...,n},B=[b_{ij}]_{i,j=1,2,...,n}\in\mathbb{R}^{n\times n}\\ a_{ij}&:=\sqrt{5}(i+j-1),b_{ij}:=\sqrt{3}(n-i+1)\end{split}$ (1)
as defined in (1).
Our library supports four matrix multiplication schemes: the naive triple-loop method, a block-based algorithm, Strassen’s algorithm, and the Ozaki scheme. In the present evaluation, Strassen’s algorithm was employed to isolate the performance contribution of BF arithmetic. The recursion in Strassen’s algorithm terminates when the submatrix size falls below $32\times 32$. The performance benefit of BF arithmetic was measured for both nonvectorized and SIMD-vectorized (AVX2 and Neon) implementations. Execution times with OpenMP parallelization using 8 and 32 threads are also reported. Because our Strassen routine does not achieve further speedup beyond 8 threads, results for the block-based algorithm (B32 Threads) are additionally included for the 32-thread case only. The numerical accuracy of all computed results falls within the ranges of 29.0–30.4 significant decimal digits for DD, 45.5–47.5 for TD, and 61.7–63.5 for QD precision.
Figure 1 shows the computation times for real square matrix multiplication at TD and QD precision. The horizontal axis represents the matrix size n=32n=3220492049, and the vertical axis represents the computation time in seconds.
Figure 1: Computation time (s) of Strassen matrix multiplication: Snapdragon (left), EPYC (right)
Notably, BF arithmetic provides no benefit—rather degrades performance—for DD precision, whereas it yields clear speedups for TD and QD precisions. For $n=2049$, the resulting speedups over the MPBLAS [13] DD- and QD-precision Rgemm are summarized in Table 1.
Table 1: Real-valued matrix multiplication time (s) for $n=2049$
MPBLAS S+SIMD BF S+SIMD 8 Threads 32 Threads B32 Threads
$D$ Snapdragon 13.6 9.7 15.4 2.5
EPYC 13.9 6.7 10.3 1.5 1.4 0.9
TD Snapdragon 61.7 39.9 8.1
EPYC 46.1 30.4 4.8 4.7 3.4
QD Snapdragon 577.5 101.4 65.6 12.3
EPYC 781.0 117.2 59.9 9.2 9.1 6.2
Table 2 reports the average speedup ratios owing to BF arithmetic for DD precision, where BF arithmetic offers no benefit, comparing nonSIMD (NS) and SIMD-vectorized implementations.
Table 2: Average speedup with branch-free (BF) arithmetic: standard / BF
Speedup $D$ TD QD
ratio NS SIMD NS SIMD NS SIMD
Snapdragon 0.6 0.7 1.2 1.4 2.0 1.4
EPYC 1.0 0.7 1.4 1.5 1.5 1.9
[…]
3.2 Benchmark Tests for Complex Matrix Multiplication
Owing to the superiority of BF arithmetic for TD and QD computations demonstrated previously, the complex matrix multiplication benchmarks employed BF-based implementations for comparison. The complex square matrix multiplication is defined using a uniform random variable $r_{u}\in[0,1]$ and a standard normal random variable $r_{n}\in\mathbb{R}$ as
$\begin{split}C&:=AB,\\ A&=[a_{ij}]_{i,j=1,2,...,n},B=[b_{ij}]_{i,j=1,2,...,n}\in\mathbb{R}^{n\times n}\\ a_{ij}&:=\exp(r_{n})(r_{u}-1/2)+\exp(r_{n})(r_{u}-1/2)\sqrt{-1}\\ b_{ij}&:=\exp(r_{n})(r_{u}-1/2)+\exp(r_{n})(r_{u}-1/2)\sqrt{-1}\end{split}$ (2)
Our complex matrix multiplication is implemented using three real matrix multiplications via the 3M method, a variant of the Karatsuba algorithm [10]. Therefore, both the Strassen and block-based algorithms perform three real matrix multiplications. The numerical accuracy, measured as the minimum of the real and imaginary parts, falls within the ranges of 23.2–24.4 significant decimal digits for DD, 39.2–40.8 for TD, and 55.8–56.5 for QD precision.
Figure 2 presents the benchmark results for complex matrix multiplication on the EPYC and Snapdragon platforms.
Figure 2: Computation time (s) of Strassen matrix multiplication: Snapdragon (left), EPYC (right)
The results demonstrate performance characteristics consistent with those observed for real matrix multiplication. In particular, extracting the computation times for $n=2049$ reveals that our implementation achieved speedups significantly exceeding those for real matrix multiplication relative to MPBLAS, as listed in Table 3.
Table 3: Complex-valued matrix multiplication time (s) for $n=2049$
Precision Platform MPBLAS S+SIMD 8 Threads 32 Threads B32 Threads
$D$ Snapdragon 72.1 29.9 7.8
EPYC 254.3 20.4 5.0 4.4 3.0
TDBF Snapdragon 123.2 24.4
EPYC 90.5 14.6 14.4 10.3
QDBF Snapdragon 2004.2 203.4 34.0
EPYC 3463.1 196.1 29.2 29.2 19.4
[…]
4 Benchmark Tests for Polynomial Evaluation and Algebraic Equation Solvers
As demonstrated in the matrix multiplication benchmarks, BF arithmetic offers no speedup for DD precision when the conventional sloppy addition and multiplication are employed. In this section, we present benchmark results for polynomial evaluation and an algebraic equation solver based on the Durand–Kerner (DK) method, both of which utilize TD and QD precision arithmetic, for which performance improvements are anticipated.
4.1 Evaluation of Real-Coefficient Polynomial Functions
For polynomial evaluation, coefficients of the degree-$n$ real-coefficient polynomial $p_{n}(x)=\sum^{n}_{i=0}a_{i}x^{i}$ are generated randomly, and the results are compared among the standard Horner’s method, Estrin’s method, and SIMD-vectorized Estrin’s method.
[…]
Figure 3 shows the computation time (in microseconds) for evaluating the real-coefficient polynomial $p_{n}(x)$ at real arguments. The upper panel corresponds to the Snapdragon platform and the lower panel to the EPYC platform; the standard arithmetic (NoBF) is compared against the branch-reduction BF arithmetic (BF). As the degree (or equivalently, the evaluation count corresponding to the problem size) increases, computation time grows, and the effects of branch elimination and vectorization become more pronounced, particularly for high-precision variants such as TD and QD, where operation count and cost are dominant. Across both architectures, a consistent trend is observed wherein BF reduces computation time relative to NoBF for TD and QD precisions.
[…]
Figure 4 shows the computation time (in microseconds) for evaluating the real-coefficient polynomial $p_{n}(x)$ at complex arguments. As each complex multiplication and addition decomposes into multiple real operations, computation times tend to be higher than those for real-argument evaluation shown in Fig. 3. By contrast, as the number of operations increases, the impact of conditional branching and normalization steps also grows; therefore, BF arithmetic should in principle remain effective for complex-argument evaluation. Although the Snapdragon results largely conform to this expectation, the EPYC results show slight discernible difference; in particular, the performance difference for TD arithmetic appears to be minimal, even when one exists.
4.2 Algebraic Equation Solver Based on the DK Method
The method under comparison is the quadratically convergent DK method, a classical simultaneous iterative solver. Denoting the approximate roots after $k$ iterations as
$\mathbf{z}_{k}=[z^{(k)}_{1}\ z^{(k)}_{2}\ ...\ z^{(k)}_{n}]^{T}\in\mathbb{C}^{n},$
the recurrence relation, expressed using the monic polynomial $q_{n}(x)=p_{n}(x)/a_{n}$, is given by
$z^{(k+1)}_{i}:=z^{(k)}_{i}-\frac{q_{n}(z^{(k)}_{i})}{\prod^{n}_{j=1,j\not=i}(z^{(k)}_{i}-z^{(k)}_{j})}$ (3)
as the quadratically convergent DK method.
The Aberth initial values are expressed by
$z^{(0)}_{i}:=-\frac{c_{n-1}}{n}+r\exp\left\{\left(\frac{2(i-1)\pi}{n}+\frac{3}{2n}\right)\mathrm{i}\right\}$ (4)
In the experiments, the radius $r$ was defined using the number of nonzero coefficients $n_{\rm nz}\leq n$ as
$r:=\max_{0\leq i\leq(n-1)}|n_{\rm nz}\,c_{i}|^{1/(n-i)}.$
The test problem was the Chebyshev integration problem, in which the polynomial coefficients, with $a_{n}=1$, are computed as follows:
[…]
where $k=1,2,\ldots,\lfloor n/2\rfloor$.
Table 4 presents the benchmark results obtained by applying the DK method with DD, TD, and QD precision arithmetic to the Chebyshev integration problem described previously.
Table 4: Benchmark test for the Chebyshev integration problem ($n=64$)
Snapdragon $D$ TD QD
NoBF 0.0528 0.174 0.59
NoBF Neon 0.0381 0.172 0.47
BF 0.14 0.383
BF Neon 0.111 0.336
EPYC $D$ TD QD
NoBF 0.0627 0.306 0.989
NoBF AVX2 0.0431 0.251 0.608
BF 0.214 0.628
BF AVX2 0.156 0.415
[…]
5 Conclusion and Future Work
In this study, we applied BF algorithms, which suppress conditional branching, to multi-component multiple-precision arithmetic (DD/TD/QD) and evaluated their performance on CPU SIMD (AVX2 on EPYC) and Arm Neon. Benchmark tests were conducted for real and complex square matrix multiplication, real- and complex-valued evaluation of real-coefficient polynomials, and the Chebyshev integration problem. The results confirm that, for high precision (particularly TD and QD), the combined use of BF conversion and SIMD vectorization is effective and reduces computation time compared with the standard implementation (NoBF). In addition, the results validate that BF conversion offers no benefit for DD precision arithmetic. These findings indicate that achieving performance improvements in multi-component arithmetic requires (i) branch elimination (BF conversion) and (ii) exploitation of instruction-level parallelism (SIMD vectorization) to be combined, with the implementation strategy selected according to the target problem, precision level, and architecture.
The following directions are identified for future work.
  1. 1.
    Adaptive implementation selection (autotuning): As the optimal execution path (NoBF/BF, SIMD width, and loop partitioning) varies with the precision level (DD/TD/QD) and problem size, we will introduce a mechanism for selecting the optimal implementation either at runtime or at build time.
  2. 2.
    Support for wider SIMD extensions and instruction sets: Extensions to AVX-512 and Arm SVE/SVE2 will be pursued, along with establishing design guidelines for data layout and reduction operations as register width increases.
  3. 3.
    Memory access optimization and enhanced parallelization: In addition to improving cache efficiency in matrix operations and polynomial evaluation, we will optimize the coordination of thread-level parallelism via OpenMP with SIMD vectorization, including NUMA placement and scheduling.
  4. 4.
    Extension to application problems: The scope of application will be expanded to include high-precision linear algebra problems (QR decomposition, eigenvalue problems, and iterative methods), special function evaluation, numerical integration, and optimization, and the effectiveness in practical settings (convergence behavior, total runtime, and energy efficiency) will be evaluated.
Through these efforts, we aim to realize the accelerated multi-component high-precision arithmetic in a more general and portable form, establishing a high-precision numerical computing infrastructure applicable to a wide range of computational environments from mobile devices to supercomputers.
Acknowledgment
This research was supported by JSPS KAKENHI Grant Number 23K11127.